# The Large Hadron electron Collider as a bridge project for CERN

### v2

@@ II The LHeC at the frontline of particle and nuclear physics
The physics program at the LHeC [1, 2, 11], based on the extended kinematic plane that it can explore (illustrated in Fig. 2), is both comprehensive and distinctive, encompassing a broad spectrum of topics. The LHeC will provide centre-of-mass energies $\sim 1.2\ (0.8)$  TeV/nucleon and integrated luminosities of 1 ab-1 (10 fb-1) in $ep$ ($eA$) collisions, respectively, with 50 fb-1 in $ep$ achievable in a first year of running. The high centre-of-mass energy, together with the large integrated luminosity that is enabled by the high-intense HL-LHC proton beam and the ERL electron beam, enables the measurement of a large variety of processes, as displayed in Fig. 3.
Figure 2: Coverage of the kinematic plane in deep inelastic lepton-proton scattering by some initial fixed target experiments, with electrons (SLAC) and muons (NMS, BCDMS), and by the $ep$ colliders: the EIC (green), HERA (yellow), the LHeC (blue) and the FCC-eh (brown). $x$ is the momentum fraction of the proton or nucleus taken by the partons involved in the cross section for a given observable, and $Q^{2}$ the hard scale of the process squared. The physics topics that the LHeC can cover are illustrated in the kinematic plane. Figure modified from [2].
Figure 3: Selected inclusive production cross sections for an electron–proton collider at the HL-LHC as a function of the electron-beam energy $E_{e}$. The proton-beam energy is set by the HL-LHC to 7 TeV, so that the $ep$ centre-of-mass energy and the cross sections depend only on $E_{e}$ of the to-be-build electron accelerator. Displayed are inclusive production cross sections for prompt $Z$ and $W^{\pm}$ bosons, the Higgs boson ($H$), single top-quark production ($t$) and top-quark pair production ($t\bar{t}$), each for neutral-current (NC) and charged-current (CC) induced channels, respectively. Furthermore, the cross sections for $HZ$ and $HW$ production are displayed, which are mainly induced by vector-boson-fusion channels. Charged-current mediated cross sections scale as  $(1+P_{e})$,  with $P_{e}$ being the longitudinal electron beam polarization which is set to $P_{e}=80\%$. For $E_{e}=50$ GeV, the total inclusive NC and CC DIS production cross sections, $6.5\cdot 10^{8}$ fb and $3.0\cdot 10^{5}$ fb, respectively, are not displayed, nor more specialized processes like $b\bar{b}$, $c\bar{c}$, jets, prompt photons, and di-boson production ($WW$, $ZZ$), or others like photoproduction or $\gamma\gamma$ induced ones. For example, $t\bar{t}$ photoproduction has a cross section of $0.7$ pb at $E_{e}=60$ GeV [12] which is more than a factor of 30 larger than NC DIS $t\bar{t}$ production displayed here (violet dashed). The right-sided axis displays the expected number of events for an integrated luminosity of $\mathcal{L}=1\,\text{ab}^{-1}$. The predictions are calculated with Sherpa3 [13] at next-to-leading order QCD.
@@ II.2 Higgs Boson Physics
The LHeC is sensitive to the six dominant Higgs decay channels: $b\bar{b},~{}W^{+}W^{-},~{}gg,~{}\tau\tau,~{}c\bar{c}$, and $ZZ$, which together account for 99.6% of the total SM Higgs decay width. The respective relative uncertainties on the expected measurements of Higgs couplings from the LHeC for the seven most abundant decay channels are summarized in Tab. 1. High precision measurements with uncertainties at the 1.9 (0.70, 1.2)level of Higgs couplings to $b$-quarks ($W$-bosons, $Z$ bosons), are achievable (the latter particularly clean in $WW\to H\to WW$ (CC) and $ZZ\to H\to ZZ$ (NC) processes), enabling reconstruction of the total SM decay width with percent-level accuracy [2]. As discussed in Sect. III, combining these measurements with HL-LHC data will further enhance precision, turning the LHeC into a Higgs high precision facility.
Coupling $\kappa$ for decay $H\to$ $b\bar{b}$ $WW$ $gg$ $\tau\tau$ $c\bar{c}$ $ZZ$ $\gamma\gamma$
@@
Table 1: Summary of uncertainties of Higgs couplings from $ep$ for the seven most abundant decay channels [2].
@@ II.3 Top-Quark Physics
Table 2: Expected limits at 95% CL on anomalous right-handed vector ($f^{1}_{R}$), left-handed tensor ($f^{2}_{L}$) and right-handed tensor ($f^{2}_{R}$) $Wtb$ couplings at the LHeC [2] and the HL-LHC [28].
@@
Figure 5: Expected sensitivities on the SM $|V_{tb}|$ ($f^{1}_{L}\equiv 1+\Delta f^{1}_{L}$) and on anomalous right-handed vector ($f^{1}_{R}$), left-handed tensor ($f^{2}_{L}$) and right-handed tensor ($f^{2}_{R}$) $Wtb$ couplings at the LHeC [25] (left), and on $|V_{ts}|$ exploring three different signal scenarios (Signal 1: $pe^{-}\to\nu_{e}\bar{t}\to\nu_{e}W^{-}\bar{b}\to\nu_{e}\ell^{-}\nu_{\ell}\bar{b}$; Signal 2: $pe^{-}\to\nu_{e}W^{-}b\to\nu_{e}\ell^{-}\nu_{\ell}b$; Signal 3: $pe^{-}\to\nu_{e}\bar{t}\to\nu_{e}W^{-}j\to\nu_{e}\ell^{-}\nu_{\ell}j$[30] (right), as a function of the integrated luminosity.
Additionally, the LHeC provides an ideal setting for precise studies of top-quark properties, such as spin and polarization, with theoretical and experimental advantages over the LHC [35]. Furthermore, the LHeC also provides access to Flavor Changing Neutral Current (FCNC) processes via $\gamma tq$ and $Ztq$ vertices ($q=u,c$[36, 37]. Processes such as $e^{-}p\to e^{-}\,W\,b+X$ can be precisely measured, improving current limits on $\gamma tu$ couplings and achieving competitiveness with HL-LHC projections.
@@ II.4 Precision SM Physics
The colour-neutral electron scattering off a coloured proton parton provides direct access to QCD phenomena, with precision enhanced by advanced vertexing, tracking, and highly granular calorimetry, as well as the unique in-situ calibration techniques in NC DIS. At the LHeC, the strong coupling constant $\alpha_{s}(M_{Z})$ can be determined with exceptional precision through scaling violations and jet measurements in the Breit frame. The precise measurement of inclusive NC and CC DIS cross sections provide precise data for the determination of the PDFs. Through higher-order corrections and the DGLAP formalism, the strong coupling can be determined from these data simultaneously with the PDF parameters, and an uncertainty of ±$\pm 0.00022$ will be achieved [2], a fourfold improvement over the current world average [38].
Inclusive jet cross sections provide further access to $\alpha_{s}(M_{Z})$. Leveraging high-resolution calorimetry, in situ calibration, and a jet-energy scale uncertainty of about 0.5%, $\alpha_{s}(M_{Z})$ can be measured with an uncertainty of ±$\pm 0.00016$ [2], a tenfold improvement over HERA results [39, 40]. These measurements, free of higher-twist and nuclear corrections, challenge lattice results and demand improved higher-order QCD predictions. Moreover, the scale dependence of $\alpha_{s}$ can be probed over a wide range, as can be seen in Fig. 6, providing a unique test of the $SU_{c}(3)$ gauge theory.
@@ II.5 Beyond the Standard Model Searches
A notable example involves heavy neutrinos (HNs). At the LHeC they can be produced via charged current interactions, with cross sections dependent on the active-sterile mixing parameter $|\theta_{e}|^{2}$. The most promising channels involve lepton flavor violation and displaced vertices, with strong prospect of discovery [55] as presented in Fig. 9 (left). The LHeC sensitivity to heavy neutrinos above 100 GeV surpasses that of $pp$, due to the lower backgrounds, or $e^{+}e^{-}$ colliders. Detection prospects are further improved using jet substructure techniques for boosted $W$ bosons from $N\to eW\to ejj$ decays, which enhances signal discrimination. With 1 ab-1 of data, the LHeC could significantly improve constraints on $|V_{eN}|^{2}$ over current LHC, $0\nu 2\beta$, and EWK limits. A recent study [56] shows that the lepton number violation signal process $e^{-}p\to\tau^{+}jjj$ can be utilized to search for heavy Majorana neutrinos. Results show that the LHeC can probe large additional regions in the parameter space formed by $|V_{\tau N}|^{2}$ and $|V_{eN}|^{2}$, with best discovery sensitivity for $m_{N}\simeq 100$ GeV. Similar conclusions are extracted in the framework of the $\nu$SMEFT [57].
@@ II.6 High-energy and high-density measurements of heavy ion collisions
Figure 10: Relative difference, in %, for the longitudinal structure function $(F_{L}^{\rm BK}-F_{L}^{\rm Rew})/F_{L}^{\rm BK}$ between the saturation predictions (labeled ”BK”) structure functions and the matched $F_{\rm L}$ for proton (left) and Au (as example of heavy nuclei, right) obtained from DGLAP fits and named ”Rew”, as a function of $x$ and $Q^{2}$. The color scale/axis goes in a linear scale from $-10\,\%$ to $10\,\%$ and in a logarithmic scale outside that range. The black dots indicate the points where results from DGLAP fits are forced to match those of saturation. Plots taken from [76].
@@ III.1 LHeC-improved PDFs at hadron colliders
Figure 11: Parton-parton luminosities as a function of rapidity ($gg$ on the left, $q\bar{q}$ on the right), at $\sqrt{s}=14$  eV (top), and as a function invariant mass at $\sqrt{s}=14$ (middle) and 100 TeV (bottom), normalized to PDF4LHC21 [15]. LHeC results at NNLO are shown with uncertainty bands together with central values of ABMP16 [16], CT18 [17], MSHT20 [18] and NNPDF4.0 [19].
@@ III.2 Impact on hadron-collider physics
Current prospects of the HL-LHC precision in Higgs and electroweak measurements are limited by strong-interaction uncertainties. A primary example of the impact of LHeC data on the physics program of the HL-LHC is given by the enhanced precision in the prediction of the Higgs boson cross section in the Standard Model. A breakdown of $gg\to H$ cross section uncertainties, taken from [81] and updated for recent progress [82, 83, 84, 85], is given in Tab. 3. The numbers in Tab. 3 show the present perturbative and PDF$+\alpha_{S}$ uncertainty; scenario ”S2” of [81] roughly corresponds to a halving of these uncertainties on the timescale of the HL-LHC. The expected impact of LHeC on PDFs and $\alpha_{S}$ is estimated following [14], with an improvement factor of 2.5–3 with respect to S2, reducing the corresponding uncertainty in the $gg\to H$ cross section to about 0.5% and making this contribution negligible with respect to the anticipated perturbative uncertainty. This opens up the possibility of a significant decrease in uncertainty when the theoretical errors are further reduced.
$\sqrt{s}$ [TeV] $\sigma_{gg\to H}$ [pb] TH uncertainty PDF+$\alpha_{S}$ uncertainty Total
@@ IV LHeC in the global context
Figure 14: Relative uncertainty in the coupling modifiers obtained in the kappa-3 framework [95, 96], for different combinations of the HL-LHC, LHeC, and other future collider datasets. ”HL-LHC(improved)” refers to the impact of reduced PDF+$\alpha_{s}$ uncertainties in the HL-LHC measurements. These data are further combined with LHeC Higgs-boson measurements, and with either FCC-ee data at $\sqrt{s}=240$ and $365$ GeV (left), or with FCC-hh and CEPC data (right). The fermion, vector, and loop modifiers are shown from top to bottom respectively. Since neither HL-LHC nor LHeC give direct access to the Higgs width, the condition $\kappa_{V}(\kappa_{W},\kappa_{Z})\leq 1$ is imposed on the fit and the uncertainty on these modifiers is defined as $1-\kappa_{V}(68\%)$. Results have been obtained with the SMEFiT framework [97, 98, 99, 100].
@@
Figure 15: Id. to Fig. 14 (right column) but showing the combination HL-LHC+FCC-hh in absence of LHeC improvements and data instead of the combination with the CEPC. The fermion, vector, and loop modifiers are shown on the top left, top right and bottom plots, respectively. Results have been obtained with the SMEFiT framework [97, 98, 99, 100].
@@ VII PERLE: A Multi-Turn, High Power Energy Recovery Linac (ERL) Facility
ERL is one of the five ESPP identified keys areas for an intensification of R&D [114, 115]. Major benefits arise from operation in ERL mode: (i) reduced RF power requirements, with the RF power needed being significantly lower as it only needs to excite the cavity and compensate for minor losses, thus leading to unprecedented intrinsic efficiency compared to conventional linacs (factor of reduction $>$15 for PERLE); (ii) minimized radioprotection, through dumping the beam at low energy, below the activation threshold, which minimizes the beam dump, thus limiting the environmental impact of the accelerator throughout its life cycle, and; (iii) excellent beam quality for experiments, as experiments can benefit from a continuously injected beam, with specifications being defined by the source, because the used beam is sent to the dump after recovering its energy.
@@ VII.1 Experimental Facility and Specifications
PERLE (Powerful ERL for Experiments, [10, 116, 117]) is a novel multi-turn, ERL facility currently under construction at IJCLab in Orsay, France. Its primary objective is to demonstrate high-current (20 mA), continuous wave (CW), multi-pass operation using superconducting cavities at 801.58 MHz. PERLE will serve as a hub for validating and exploring a wide range of accelerator phenomena in an unprecedented operational power regime, contributing to the development of ERL technology for future high-energy and high-intensity machines. It is specifically designed to validate choices and provide experience with operation for an ERL envisioned in the design of the LHeC and the FCC-eh. In addition, PERLE will provide a high-energy ($>$100 keV) X-ray source through inverse Compton process thanks to a high finesse Fabry-Perot optical cavity. Moreover, PERLE could host unique experiments in nuclear physics, being the first ERL to study electron-nucleus interactions with radioactive nuclei.
@@ VIII.2 LHeC as a Detector Development Stepping-Stone towards Higgs factories
The need for precision electromagnetic calorimetry at Higgs Factories (a few %$\%/\surd E$) is mirrored in DIS experiments where they play a crucial role in the pivotal task of scattered electron reconstruction. The use of silicon detectors in forward and backward calorimetry has become more common recently, as is notably being deployed in CMS [101]. Further forward, the silicon pixel-based ALICE FOCAL will come on-line on the HL-LHC timescale and encourages the use of similar solutions even in beam-line instrumentation, as has been considered for the zero degree calorimeter at ePIC and in the baseline LHeC design. The current LHeC design already widely deploys silicon in sampling calorimetry for the end-cap regions, offering a clear testing ground to Higgs factories. At ePIC, the deployment of silicon layers extends to the barrel region, where 4 astropix layers are interleaved with a basic lead-scintillating fibre design to provide precision pointing capabilities in the electromagnetic calorimetry. Whereas the current LHeC barrel electromagnetic calorimeter baseline design, see Sec. V.1, is an evolution of the liquid argon technology in use by ATLAS, there is scope to adjust, particularly if there is a motivation from future field-wide needs.

### v3

@@ II The LHeC at the frontline of particle and nuclear physics
The physics program at the LHeC [1, 2, 11], based on the extended kinematic plane that it can explore (illustrated in Fig. 2), is both comprehensive and distinctive, encompassing a broad spectrum of topics. The LHeC will provide centre-of-mass energies $\sim 1.2\ (0.8)$  TeV/nucleon and integrated luminosities of 1 ab-1 (10 fb-1) in epep (eAeA) collisions, respectively, with 50 fb-1 in $ep$ achievable in a first year of running. The high centre-of-mass energy, together with the large integrated luminosity that is enabled by the high-intense HL-LHC proton beam and the ERL electron beam, enables the measurement of a large variety of processes, as displayed in Fig. 3.
Figure 2: Coverage of the kinematic plane in deep inelastic lepton-proton scattering by some initial fixed target experiments, with electrons (SLAC) and muons (NMS, BCDMS), and by the epep colliders: the EIC (green), HERA (yellow), the LHeC (blue) and the FCC-eh (brown). $x$ is the momentum fraction of the proton or nucleus taken by the partons involved in the cross section for a given observable, and $Q^{2}$ the hard scale of the process squared. The physics topics that the LHeC can cover are illustrated in the kinematic plane. Figure modified from [2].
Figure 3: Selected inclusive production cross sections for an electron–proton collider at the HL-LHC as a function of the electron-beam energy $E_{e}$. The proton-beam energy is set by the HL-LHC to 7 TeV, so that the epep centre-of-mass energy and the cross sections depend only on $E_{e}$ of the to-be-build electron accelerator. Displayed are inclusive production cross sections for prompt $Z$ and $W^{\pm}$ bosons, the Higgs boson ($H$), single top-quark production ($t$) and top-quark pair production ($t\bar{t}$), each for neutral-current (NC) and charged-current (CC) induced channels, respectively. Furthermore, the cross sections for $HZ$ and HWHW production are displayed, which are mainly induced by vector-boson-fusion channels. Charged-current mediated cross sections scale as  $(1+P_{e})$,  with $P_{e}$ being the longitudinal electron beam polarization which is set to $P_{e}=80\%$. For $E_{e}=50$ GeV, the total inclusive NC and CC DIS production cross sections, $6.5\cdot 10^{8}$ fb and $3.0\cdot 10^{5}$ fb, respectively, are not displayed, nor more specialized processes like $b\bar{b}$, $c\bar{c}$, jets, prompt photons, and di-boson production (WWWW, ZZZZ), or others like photoproduction or $\gamma\gamma$ induced ones. For example, $t\bar{t}$ photoproduction has a cross section of $0.7$ pb at $E_{e}=60$ GeV [12] which is more than a factor of 30 larger than NC DIS $t\bar{t}$ production displayed here (violet dashed). The right-sided axis displays the expected number of events for an integrated luminosity of $\mathcal{L}=1\,\text{ab}^{-1}$. The predictions are calculated with Sherpa3 [13] at next-to-leading order QCD.
@@ II.2 Higgs Boson Physics
The LHeC is sensitive to the six dominant Higgs decay channels: $b\bar{b},~W^{+}W^{-},~gg,~\tau\tau,~c\bar{c}$, and $ZZ$, which together account for 99.6% of the total SM Higgs decay width. The respective relative uncertainties on the expected measurements of Higgs couplings from the LHeC for the seven most abundant decay channels are summarized in Tab. 1. High precision measurements with uncertainties at the 1.9 (0.70, 1.2)% level of Higgs couplings to $b$-quarks ($W$-bosons, $Z$ bosons), are achievable (the latter particularly clean in $WW\to H\to WW$ (CC) and $ZZ\to H\to ZZ$ (NC) processes), enabling reconstruction of the total SM decay width with percent-level accuracy [2]. As discussed in Sect. III, combining these measurements with HL-LHC data will further enhance precision, turning the LHeC into a Higgs high precision facility.
Coupling $\kappa$ for decay $H\to$ $b\bar{b}$ WWWW gggg $\tau\tau$ $c\bar{c}$ ZZZZ $\gamma\gamma$
@@
Table 1: Summary of uncertainties of Higgs couplings from epep for the seven most abundant decay channels [2].
@@ II.3 Top-Quark Physics
Table 2: Expected limits at 95% CL on anomalous right-handed vector ($f^{1}_{R}$), left-handed tensor ($f^{2}_{L}$) and right-handed tensor ($f^{2}_{R}$) WtbWtb couplings at the LHeC [2] and the HL-LHC [28].
@@
Figure 5: Expected sensitivities on the SM $|V_{tb}|$ ($f^{1}_{L}\equiv 1+\Delta f^{1}_{L}$) and on anomalous right-handed vector ($f^{1}_{R}$), left-handed tensor ($f^{2}_{L}$) and right-handed tensor ($f^{2}_{R}$) WtbWtb couplings at the LHeC [25] (left), and on $|V_{ts}|$ exploring three different signal scenarios (Signal 1: $pe^{-}\to\nu_{e}\bar{t}\to\nu_{e}W^{-}\bar{b}\to\nu_{e}\ell^{-}\nu_{\ell}\bar{b}$; Signal 2: $pe^{-}\to\nu_{e}W^{-}b\to\nu_{e}\ell^{-}\nu_{\ell}b$; Signal 3: $pe^{-}\to\nu_{e}\bar{t}\to\nu_{e}W^{-}j\to\nu_{e}\ell^{-}\nu_{\ell}j$[30] (right), as a function of the integrated luminosity.
Additionally, the LHeC provides an ideal setting for precise studies of top-quark properties, such as spin and polarization, with theoretical and experimental advantages over the LHC [35]. Furthermore, the LHeC also provides access to Flavor Changing Neutral Current (FCNC) processes via $\gamma tq$ and $Ztq$ vertices (q=u,cq=u,c[36, 37]. Processes such as $e^{-}p\to e^{-}\,W\,b+X$ can be precisely measured, improving current limits on $\gamma tu$ couplings and achieving competitiveness with HL-LHC projections.
@@ II.4 Precision SM Physics
The colour-neutral electron scattering off a coloured proton parton provides direct access to QCD phenomena, with precision enhanced by advanced vertexing, tracking, and highly granular calorimetry, as well as the unique in-situ calibration techniques in NC DIS. At the LHeC, the strong coupling constant $\alpha_{s}(M_{Z})$ can be determined with exceptional precision through scaling violations and jet measurements in the Breit frame. The precise measurement of inclusive NC and CC DIS cross sections provide precise data for the determination of the PDFs. Through higher-order corrections and the DGLAP formalism, the strong coupling can be determined from these data simultaneously with the PDF parameters, and an uncertainty of $\pm 0.00022$ will be achieved [2], a fourfold improvement over the current world average [38].
Inclusive jet cross sections provide further access to $\alpha_{s}(M_{Z})$. Leveraging high-resolution calorimetry, in situ calibration, and a jet-energy scale uncertainty of about 0.5%, $\alpha_{s}(M_{Z})$ can be measured with an uncertainty of $\pm 0.00016$ [2], a tenfold improvement over HERA results [39, 40]. These measurements, free of higher-twist and nuclear corrections, challenge lattice results and demand improved higher-order QCD predictions. Moreover, the scale dependence of $\alpha_{s}$ can be probed over a wide range, as can be seen in Fig. 6, providing a unique test of the $SU_{c}(3)$ gauge theory.
@@ II.5 Beyond the Standard Model Searches
[…]
@@ II.6 High-energy and high-density measurements of heavy ion collisions
Figure 10: Relative difference, in %, for the longitudinal structure function $(F_{L}^{\rm BK}-F_{L}^{\rm Rew})/F_{L}^{\rm BK}$ between the saturation predictions (labeled ”BK”) structure functions and the matched $F_{\rm L}$ for proton (left) and Au (as example of heavy nuclei, right) obtained from DGLAP fits and named ”Rew”, as a function of $x$ and $Q^{2}$. The color scale/axis goes in a linear scale from $10\,\%$ to $10\,\%$ and in a logarithmic scale outside that range. The black dots indicate the points where results from DGLAP fits are forced to match those of saturation. Plots taken from [76].
@@ III.1 LHeC-improved PDFs at hadron colliders
Figure 11: Parton-parton luminosities as a function of rapidity (gggg on the left, $q\bar{q}$ on the right), at $\sqrt{s}=14$  eV (top), and as a function invariant mass at $\sqrt{s}=14$ (middle) and 100 TeV (bottom), normalized to PDF4LHC21 [15]. LHeC results at NNLO are shown with uncertainty bands together with central values of ABMP16 [16], CT18 [17], MSHT20 [18] and NNPDF4.0 [19].
@@ III.2 Impact on hadron-collider physics
[…]
@@ IV LHeC in the global context
[…]
@@
Figure 15: Id. to Fig. 14 (right column) but showing the combination HL-LHC+FCC-$h$ in absence of LHeC improvements and data instead of the combination with the CEPC. The fermion, vector, and loop modifiers are shown on the top left, top right and bottom plots, respectively. Results have been obtained with the SMEFiT framework [97, 98, 99, 100].
@@ VII PERLE: A Multi-Turn, High Power Energy Recovery Linac (ERL) Facility
ERL is one of the five ESPP identified keys areas for an intensification of R&D [114, 115]. Major benefits arise from operation in ERL mode: (i) reduced RF power requirements, with the RF power needed being significantly lower as it only needs to excite the cavity and compensate for minor losses, thus leading to unprecedented intrinsic efficiency compared to conventional linacs (factor of reduction >>15 for PERLE); (ii) minimized radioprotection, through dumping the beam at low energy, below the activation threshold, which minimizes the beam dump, thus limiting the environmental impact of the accelerator throughout its life cycle, and; (iii) excellent beam quality for experiments, as experiments can benefit from a continuously injected beam, with specifications being defined by the source, because the used beam is sent to the dump after recovering its energy.
@@ VII.1 Experimental Facility and Specifications
PERLE (Powerful ERL for Experiments, [10, 116, 117]) is a novel multi-turn, ERL facility currently under construction at IJCLab in Orsay, France. Its primary objective is to demonstrate high-current (20 mA), continuous wave (CW), multi-pass operation using superconducting cavities at 801.58 MHz. PERLE will serve as a hub for validating and exploring a wide range of accelerator phenomena in an unprecedented operational power regime, contributing to the development of ERL technology for future high-energy and high-intensity machines. It is specifically designed to validate choices and provide experience with operation for an ERL envisioned in the design of the LHeC and the FCC-eh. In addition, PERLE will provide a high-energy (>>100 keV) X-ray source through inverse Compton process thanks to a high finesse Fabry-Perot optical cavity. Moreover, PERLE could host unique experiments in nuclear physics, being the first ERL to study electron-nucleus interactions with radioactive nuclei.
@@ VIII.2 LHeC as a Detector Development Stepping-Stone towards Higgs factories
The need for precision electromagnetic calorimetry at Higgs Factories (a few $\%/\surd E$) is mirrored in DIS experiments where they play a crucial role in the pivotal task of scattered electron reconstruction. The use of silicon detectors in forward and backward calorimetry has become more common recently, as is notably being deployed in CMS [101]. Further forward, the silicon pixel-based ALICE FOCAL will come on-line on the HL-LHC timescale and encourages the use of similar solutions even in beam-line instrumentation, as has been considered for the zero degree calorimeter at ePIC and in the baseline LHeC design. The current LHeC design already widely deploys silicon in sampling calorimetry for the end-cap regions, offering a clear testing ground to Higgs factories. At ePIC, the deployment of silicon layers extends to the barrel region, where 4 astropix layers are interleaved with a basic lead-scintillating fibre design to provide precision pointing capabilities in the electromagnetic calorimetry. Whereas the current LHeC barrel electromagnetic calorimeter baseline design, see Sec. V.1, is an evolution of the liquid argon technology in use by ATLAS, there is scope to adjust, particularly if there is a motivation from future field-wide needs.